# Stabilizer Entropy and entanglement complexity in the Sachdev-Ye-Kitaev model

### v2

@@ Abstract
The Sachdev-Ye-Kitaev (SYK) model is of paramount importance for the understanding of both strange metals and a microscopic theory of two-dimensional gravity. We study the interplay between Stabilizer Rényi Entropy (SRE) and entanglement entropy in both the ground state and highly excited states of the SYK4+SYK2 model interpolating the highly chaotic four-body interactions model with the integrable two-body interactions one. The interplay between these quantities is assessed also through universal statistics of the entanglement spectrum and its anti-flatness. We find that SYK4 is indeed characterized by a complex pattern of both entanglement and non-stabilizer resources while SYK2 is non-universal and not complex. We discuss the fragility and robustness of these features depending on the interpolation parameter.
@@
Introduction.— The Sachdev-Ye-Kitaev (SYK) model  [73, 44] describes the behavior of strongly correlated fermions in strange metals [21]. In recent years, the model has gained new interest from the high-energy community due to its holographic duality  [43, 74, 82]. In the infra-red limit and large number of degrees of freedom, the model acquires conformal symmetry. The leading corrections to the out-of-time-ordered four-point correlation functions exhibit exponential growth over time. The rate of this growth reaches the universal upper bound established in [56], see [83, 34]. The low-energy sector of the SYK4 model is governed by an emergent reparametrization symmetry, described by Jackiw-Teitelboim (JT) gravity [82], providing insight into the holographic description of black holes and their thermodynamic properties. In quantum thermodynamics, the SYK model has been studied for its potential to show quantum advantage as a quantum battery [12, 40, 4], with a super extensive charging power [71].
The interpolation between the four-body (SYK4) and the two-body (SYK2) model [54, 30] provides a connection to a model of strongly correlated electrons. The two theories are very different: the SYK4 is chaotic, with an exponential density of states at low energy, while the SYK2 is integrable, with a polynomially vanishing gap. Remarkably, both models exhibit a volume law for entanglement [52]. One might wonder what makes these two regimes so different. Entanglement has been deeply investigated to better understand both the chaotic nature of SYK models [38, 29, 52] and their connection to holographic theory and complex quantum systems [91, 24, 21] but the transition to SYK2 shows that entanglement alone is not enough.
Figure 1: Schematic representation of the SYK4+SYK2 model $H(g)$ (see Eq. (5)). The left side ($g=0$) corresponds to SYK4 featuring the Wigner-Dyson entanglement spectrum statistics (ESS) in the ground state and near-universal capacity of entanglement $C_{E}$. On the right side ($g=1$) the SYK2 model is shown, exhibiting Poisson ESS and non-universal $C_{E}$. In the ground state, these features are fragile for SYK4 and robust for SYK2. Conversely, in high energy states the SYK4 properties extend to every value of $g\neq 1$ making the SYK2 features the fragile ones.
@@
The second layer of quantum complexity is given by non-stabilizerness [85], a crucial resource in the context of universal quantum computation, error correction [77, 11, 6, 45, 32] and quantum simulation [78, 33, 41, 13]. Recently, non-stabilizerness has risen to prominence due to finding a unique computable monotone for pure states, the Stabilizer Rényi Entropy (SRE) [51]. In the context of high-energy physics, both SRE and entanglement are studied for heavy nuclei simulations  [70] and neutrino physics [10, 20]. Moreover, the delocalization due to the entanglement of non-stablizerness resources has recently been connected to the holographic dual of back-reaction in the context of AdS-CFT [14], features of CFT [65, 86], and the harvesting of quantum resources from the vacuum of a quantum field [60, 18]. When entanglement delocalizes and scrambles non-stabilizer resources, this gives rise to universal behavior of out-of-time-order correlation functions, entanglement fluctuations [50, 64] and the onset of chaotic behavior in quantum many-body systems [87, 92, 72] in the ETH-MBL transition [88, 69, 68, 58]. On the other hand, the fine structure of entanglement reveled by the statistics of entanglement gaps (ESS) [19, 76, 92] reaches the complex universal patterns of random matrix theory thanks to non-stabilizerness.
@@
The main goal of this Letter is to study the SYK4+SYK2 model beyond perturbative limits under the lens of the emergence of quantum complex behavior from interplay between entanglement and SRE. We find that SYK4 exhibits a complex interplay of entanglement and SRE both in the ground state and in the high-energy eigenstates (middle of the spectrum), while SYK2 shows patterns of this interplay that are typical of integrable or non-chaotic models [48]. This is revealed by the adherence of SYK4 to Haar-like behavior for entanglement entropy, entanglement eigenvalues, and gaps statistics, capacity of entanglement, higher values of SRE, especially in its non-local features. On the other hand, SYK2 shows a lower level of SRE, and non-universal statistics in all the figures of merit mentioned above.
An important question is how robust are these behaviors in the SYK4+SYK2 model. The universal properties of SYK4 are fragile in the ground state: for every $g>0$ the system falls into the SYK2 class. On the other hand, high-energy states show a reversed behavior. The SYK4 phase extends to every value of $g<1$. These findings enrich those of [54] based on Green’s functions. In addition, we also show that adherence of the energy spectrum to Wigner-Dyson universal behavior is fragile for SYK4+SYK2. Finally, we show that the behavior of SRE in the ground state is capable of classifying the $8-$fold way of the symmetries of SYK4. A similar conclusion is also shown for the energy gap above the ground state. In order to argue about the robustness of these features, we apply quantum-information theoretic tools like the Kullback-Leibler divergence fidelity.
@@
The model.— The most general form of SYK models [42, 43] considers a q-body all-to-all interaction between $N$ Majorana fermionic modes. The Hamiltonian in terms of Majorana operators reads
$H_{q}=(i)^{q/2}\sum\limits_{1\leq i_{1}<\ldots<i_{q}\leq N}J_{i_{1}i_{2}\ldots i_{q}}\chi_{i_{1}}\chi_{i_{2}}\ldots\chi_{i_{q}}$ (1)
@@
with $q$ an even integer number. The disorder in the model is due to the couplings $J_{i_{1},i_{2},\dots,i_{q}}$ which are identical, independent distributed $\rm(i.i.d.)$ gaussian variables with vanishing mean and variance
$\overline{J_{i_{1},i_{2},\dots,i_{q}}}=0\;;\;\;\;\overline{J_{i_{1},i_{2},\dots,i_{q}}^{2}}=\frac{(q-1)!J}{N^{q-1}}\;.$ (2)
@@
For $q>2$, the system exhibits quantum chaos [46, 66], as evidenced by several established indicators. One key probe is level repulsion in the energy spectra, characterized by the statistical distribution of energy level spacings [2, 27, 28, 93]. This indicates that the eigenstates of the Hamiltonian are highly delocalized and correlated, a hallmark of quantum chaos. The system is defined on a complete graph, so the interactions are strongly non-local. Using the Jordan-Wigner transformations it is possible to map the Majorana operators into Pauli spin strings, used in the numerical simulations [8]. Notice that each spin operator can be expressed in terms of two Majorana operators. Therefore, the total number of Majorana operators is double the number of Paulis. The simplest version of SYK models is the SYK2
$H_{2}=i\sum\limits_{1\leq i<j\leq N}J_{i,j}\chi_{i}\chi_{j}\;\ .$ (3)
@@
Since the interactions between fermions are considered in pairs, SYK2 represents the free fermions point of the theory. This results in Gaussian statistics for its spectral properties, meaning that it does not show the same level of randomness and complexity found in chaotic systems [74, 22]. The SYK2 model serves as a simple example of a disordered fermionic system and is analytically solvable [52]. The four-body interaction model, called SYK4, is
$H_{4}=-\sum\limits_{1\leq i<j<k<l\leq N}J_{i,j,k,l}\chi_{i}\chi_{j}\chi_{k}\chi_{l}\;\;\ .$ (4)
@@
This model satisfies the usual probes of quantum chaos [66, 46], i.e. the Hamiltonian exhibits level repulsion in its spectrum. The SYK4+SYK2 model is defined as [54]
$H_{g}:=(1-g)H_{4}+gH_{2},$ (5)
with $g\in[0,1]$. In the following, we compute by exact diagonalization the energy spectrum of $H_{g}$ together with the exact ground state (GS) and middle-spectrum eigenstate (MS) for a number of realization $M$ of the disorder. The eigenvalues of the reduced density operator (RDM) to half-system $\rho_{R}$ will be denoted by $\{\lambda_{k}\}$ and are sorted in ascending order.
Entanglement.— To quantify the bipartite entanglement of the interpolated model, we focus on Rényi entropies defined as
@@
The von Neumann entanglement entropy $S(\rho_{\mathrm{R}})_{1}=-\mathrm{Tr}(\rho_{\mathrm{R}}\log(\rho_{\mathrm{R}}))$ corresponds to the limit $\alpha\to 1^{+}$.
Figure 2: Averaged bipartite entanglement in the GS (left panel) and MS (right panel) of the SYK4+SYK2 model $H_{g}$ as a function of $g$. Shaded areas represent the standard deviation across $M$ realizations of the Hamiltonian. We define $\Delta S_{1}:=|(S_{1}-S_{1}^{\rm Haar})/S_{1}^{\rm Haar}|$ the relative gap with the Haar (Page) value. The inset shows the finite size scaling of the relative gap $\Delta S_{1}$ for $g=0,1$.
For random quadratic Hamiltonians, a class of models to which the SYK2 Hamiltonian belongs [52], the average ground-state entanglement entropy was derived in closed form in [55], given by
$\mathcal{S}_{1}^{\mathrm{SYK2}}\left(R,f\right)=\mathcal{K}(f)\ln{(2)}R,$ (7)
@@
Here, $f=R/N^{\rm spin}$ represents the ratio between the subsystem size $R$ and the total system size in terms of the number of qubits. The entanglement entropy scaling of the SYK2 ground state follows a volume law (extensive scaling with subsystem size) but with a coefficient dependent on the ratio $f$, distinguishing it from the fully quantum chaotic regime in the thermodynamic limit.
Efforts to extend these insights to the SYK4 model include works such as Refs. [37, 90, 38]. On the other hand, the maximally chaotic random pure state, sampled uniformly according to the Haar measure, exhibits the Page value for entanglement entropy, given by
@@
In Fig. 2, we present the results for the rescaled entanglement entropy in the GS and MS of $H_{g}$ of the interpolated SYK Hamiltonian, with $f=1/2$. The averaging is performed over different partitions of the system (denoted with an overline $\overline{S}$)[1] and multiple disorder realizations (denoted with $\langle\rangle_{M}$) of the model for each value of the interpolation parameter $g$. Details of the realization and disorder statistics for various system sizes are provided in [39]. We see that for $g=0$ (SYK4) the GS is closer than $g=1$ (SYK2) to the Haar value, although none of them reaches the universal value. On the contrary, for MS, the relative gap $\Delta S_{1}$ shows a perfect adherence of SYK4 with Haar value unlike SYK2 as it is shown by the finite-size scaling in the inset. The behavior with $g$ suggests that the MS universal properties are valid for all $g\neq 1$, making SYK4 robust from this point of view. A similar analysis has been conducted for SYK4 in  [38].
Figure 3: Renormalized RDM eigenvalues of half-subsystem. The reference value is the Marchenko-Pastur (M-P) distribution $\eta^{\rm Haar}(x)$. We define $\eta_{k}=k/d$ and $x_{k}=(1/2)\sqrt{\lambda_{k}d}$ where $d=2^{N/4}$ for $f=1/2$.
A finer probe into the structure of entanglement is given by the full distribution of the eigenvalues of the RDM. The reference Haar value is related to the Marchenko-Pastur (M-P) distribution as the limiting distribution of eigenvalues of Wishart matrices and reads $\eta^{\rm Haar}(x)=1-\frac{2}{\pi}\left(x\sqrt{1-x^{2}}+\arcsin{x}\right)$; see [87, 57]. To distinguish two probability distributions we employ the Kullback-Leibler divergence $D_{KL}(p||q):=\sum_{i}p_{i}(\log p_{i}-\log q_{i})$. The results in the top row of Fig. 3 show the fidelity $D_{KL}(\eta_{g}|\eta_{g+\epsilon})$ between two distributions of the eigenvalues of the RDM for two nearby values $g,g+\epsilon$ of the interpolation parameter. This quantity can serve as a probe of a sharp transition associated with an observable consisting of a probability distribution. We see that the GS and MS behave in symmetric and opposite ways. The structure of the eigenvalues of the RDM of the GS shows a sharp transition at $g>0$ and then smooths out. On the other hand, for the MS, $g=1$ is fragile and the eigenvalues of the RDM of highly excited states are smoothly varying all the way as long as SYK4 interactions are different from zero. The second row of Fig. 3 shows the same statistical distance between the state for the value $g$ and the reference Haar value. While SYK4 is converging to the Haar value, SYK2 shoots away. The robustness of the two phases for GS and MS respectively is confirmed. In other words, for every value of $g\neq 1$, the eigenvalues of the RDM of highly excited states obey the universal Haar behavior.
Figure 4: ESS of the GS RDM eigenvalues of the $H_{g}$ model for different values of $g$. We superimpose the analytical curves for the Wigner-Dyson (dashed) and Poisson (blue continuous) distribution for comparison. System size $N=22$ and number of realizations is $M=100$. The number of bins used for the histogram is 100. The colored numbers represent the KL divergence of data against known distributions [39]

### v3

@@ Abstract
The Sachdev-Ye-Kitaev (SYK) model is of paramount importance for the understanding of both strange metals and a microscopic theory of two-dimensional gravity. We study the interplay between Stabilizer Rényi Entropy (SRE) and entanglement entropy in both the ground state and highly excited states of the SYK-4+SYK-2 model, interpolating the highly chaotic four-body interactions model with the integrable two-body interactions one. The interplay between these quantities is also assessed through universal statistics of the entanglement spectrum and its anti-flatness. We find that SYK-4 is indeed characterized by a complex pattern of both entanglement and non-stabilizer resources, while SYK-2 is non-universal and not complex. We discuss the fragility and robustness of these features depending on the interpolation parameter.
I Introduction
@@
The Sachdev-Ye-Kitaev (SYK) model describes the behavior of strongly correlated fermions in strange metals, and has received significant attention from the high-energy community due to its profound connection with black hole physics [1, 2, 3]. The low-energy sector of this model is described by Jackiw-Teitelboim (JT) gravity, providing insight into the holographic description of black holes and their thermodynamic properties [4, 5]. In the SYK model, the out-of-time-ordered correlators exhibit exponential growth over time with a growth rate that reaches the universal upper bound established in [6] and expected in a theory of gravity [7, 8]. Beyond this exciting connection to black hole physics, the SYK model has also been studied for its potential to be used as a quantum battery [9, 10, 11, 12], exhibiting super extensive charging power [13].
The most general variant of the model, the SYK-q, includes q-body interactions between Majorana modes, and the interpolation from the four-body SYK-4 to the two-body SYK-2 model offers a physically motivated route to understanding strongly correlated electron systems [14, 15]. The two theories are different: the SYK-4 is chaotic, with an exponential density of states at low energy, while the SYK-2 is a free random theory, with a polynomially vanishing gap [3, 4]. Remarkably, both models exhibit a volume law for entanglement [16, 17, 18, 19], prompting the question of what else sets these two models apart.
@@
In quantum information theory, it is known that beyond entanglement, the next layer of quantum complexity is captured by non-stabilizerness [20], a fundamental resource for universal quantum computation, quantum error correction [21, 22, 23, 24, 25], and quantum simulation [26, 27, 28, 29]. Recently, non-stabilizerness has risen to prominence due to finding a unique computable monotone for pure states, the Stabilizer Rényi Entropy (SRE) [30]. In the context of high-energy physics, both SRE and entanglement are studied for heavy nuclei simulations [31] and neutrino physics [32, 33]. Moreover, the delocalization due to the entanglement of non-stablizerness resources has recently been connected to the holographic dual of back-reaction in the context of AdS-CFT [34], features of CFT [35, 36], and the harvesting of quantum resources from the vacuum of a quantum field [37, 38]. When entanglement delocalizes and scrambles non-stabilizer resources, it gives rise to universal behavior (expected from states sampled according to the Haar measure) of out-of-time-ordered correlation functions [30], entanglement fluctuations [39, 40] and the onset of chaotic behavior in quantum many-body systems [41, 42, 43] in the ETH-MBL transition [44, 45, 46, 47]. On the other hand, the fine structure of entanglement revealed by the statistics of entanglement gaps (ESS) [48, 49, 42] reaches the complex universal patterns of random matrix theory thanks to non-stabilizerness.
@@
The main goal of this work is to study the SYK-4+SYK-2 model beyond the perturbative regime under the lens of the emergence of quantum complex behavior resulting from the interplay between entanglement and SRE. We find that SYK-4 exhibits a non-trivial interplay of entanglement and SRE both in the ground state and in the high-energy eigenstates (middle-of-the-spectrum states), while SYK-2 shows features that are typical of integrable or non-chaotic models [50, 51]. This is revealed by the adherence of SYK-4 to Haar-like behavior for entanglement entropy, reduced density matrix eigenvalues, and gaps statistics, capacity of entanglement, higher values of SRE, especially in its non-local features (see Fig. 1). In contrast, SYK-2 exhibits lower SRE and displays non-universal behavior across all the aforementioned figures of merit.
We address the robustness of universal features in the interpolated model, where the parameter $g$ controls the relative contribution of the SYK-4 and SYK-2 components. As defined in Eq. (5), the model reduces to SYK-4 when $g=0$, while only the SYK-2 term remains at $g=1$. Our results reveal that the universal properties of SYK-4 are highly fragile in the ground state: even for arbitrarily small $g>0$, the system transitions into the SYK-2 class. At high energies, however, this behavior is reversed—the SYK-4 phase remains robust and persists across the entire range $g<1$.
To further probe this transition, we analyze the entanglement spectrum and find that its adherence to Wigner-Dyson statistics is extremely sensitive to the interpolation, breaking down for both ground and high-energy states. To quantify this fragility, we employ tools inspired by quantum information theory, including the Kullback-Leibler (KL) divergence and a novel similarity measure we introduce, termed KL fidelity. These insights not only complement but also extend the findings of Ref. [14], which were based on Green’s function analysis. More specifically, we find that the critical interpolation parameter exhibits a power-law scaling, $g_{c}\sim N^{-3/4}$, and therefore vanishes in the thermodynamic limit. Thus, as $N\to\infty$, this shows quantitatively how the SYK-4 character of the ground state becomes unstable for arbitrarily small quadratic perturbations. Finally, we show that the behavior of SRE in the eigenstates can distinguish among the 888-fold symmetry classes of SYK-4. This is remarkable, as such classification is typically associated with properties of the full Hamiltonian and its eigenspectrum, rather than individual eigenstates.
The paper is structured as follows. We begin by introducing the SYK-q model and its interpolating version in Section II, detailing the numerical methods used for our analysis. We then study the behavior of entanglement and its spectrum across the interpolation in Sections III.1 and III.2, focusing on differences between ground and highly excited states. Next, we explore entanglement spectrum statistics in Section III.3 and the role of non-stabilizer resources, quantified by the Stabilizer Rényi Entropy in Section III.4. Finally, we investigate measures of anti-flatness in Section III.5 and discuss how they capture the non-local structure of entanglement and magic in the model. In Section IV, we provide a comprehensive summary of the findings this work and mention future perspectives.
Figure 1: Schematic representation of the interpolated SYK-4 + SYK-2 model defined in Eq. (5). At $g=0$, the model reduces to SYK-4, characterized in the ground state by Wigner-Dyson entanglement spectrum statistics (ESS) and a near-universal (Haar-like) average capacity of entanglement, $\overline{C_{E}}$. At $g=1$, the SYK-2 limit exhibits Poisson ESS and non-universal $\overline{C_{E}}$. In the ground state, SYK-4 features are fragile under interpolation, whereas SYK-2 features remain robust. In contrast, SYK-4 behavior persists at high energies for all $g\neq 1$, rendering SYK-2 traits fragile. The capacity of entanglement serves as a probe of non-local nonstabilizerness in the eigenstates of the Hamiltonian.
II The model
@@
The most general form of SYK models considers a q-body all-to-all interaction between $N$ Majorana fermionic modes with the Hamiltonian
$H^{\mathrm{SYK}-q}=(i)^{q/2}\sum\limits_{1\leq i_{1}<\ldots<i_{q}\leq N}J_{i_{1}i_{2}\ldots i_{q}}\chi_{i_{1}}\chi_{i_{2}}\ldots\chi_{i_{q}},$ (1)
@@
with $q$ an even integer number [1, 2, 3]. The disorder in the model is due to the couplings $J_{i_{1},i_{2},\dots,i_{q}}$ which are identical, independent distributed $\rm(i.i.d.)$ gaussian variables with vanishing mean and variance
$\overline{J_{i_{1},i_{2},\dots,i_{q}}}=0\;,\;\;\;\overline{J_{i_{1},i_{2},\dots,i_{q}}^{2}}=\frac{(q-1)!J}{N^{q-1}}\;.$ (2)
@@
For $q>2$, the system exhibits quantum chaos [52, 15, 53], as evidenced by several established indicators. One key probe is level repulsion in the energy spectra, characterized by the statistical distribution of energy level spacings [54, 55, 56, 57]. This indicates that the eigenstates of the Hamiltonian are highly delocalized and correlated, a hallmark of quantum chaos. The SYK model is defined on a complete graph, so the interactions are strongly non-local. Using the Jordan-Wigner transformations, it is possible to map the Majorana operators into Pauli spin strings, used in the numerical simulations [58]. Notice that each spin operator can be expressed in terms of two Majorana operators. Therefore, the total number of Majorana operators is double the number of Paulis. The simplest version of the SYK model is
$H^{\mathrm{SYK}-2}=i\!\!\!\!\sum\limits_{1\leq i<j\leq N}\!\!\!\!\!J_{i,j}\chi_{i}\chi_{j}\;.$ (3)
@@
Since the interactions between fermions are considered in pairs (see Fig. 1), SYK-2 represents the free fermions point of the theory. This results in Gaussian statistics for its spectral properties [4, 50, 51], meaning that it does not show the same level of randomness and complexity found in many-body quantum chaotic systems, and makes it an example of a disordered fermionic system and which is analytically tractable [16]. The four-body interaction model reads
$H^{\mathrm{SYK}-4}=-\!\!\!\!\!\!\!\!\!\!\sum\limits_{1\leq i<j<k<l\leq N}\!\!\!\!\!\!\!\!\!\!J_{i,j,k,l}\chi_{i}\chi_{j}\chi_{k}\chi_{l}\;.$ (4)
@@
The spectrum of the SYK-4 model exhibits clear signatures of quantum chaos, most notably level repulsion in the Hamiltonian eigenvalues [52, 15, 53, 59]. We define the interpolated the SYK-4 + SYK-2 model, as
$H(g):=(1-g)H^{\mathrm{SYK}-4}+gH^{\mathrm{SYK}-2},$ (5)
with $g\in[0,1]$ [14]. In the following, we study the interpolated model using exact numerical diagonalization, focusing on the ground state (GS) and a middle-spectrum eigenstate (MS) across many disorder realizations $M$.
III Results
III.1 Entanglement
To quantify the bipartite entanglement, we focus on Rényi entropies defined as
@@ III.1 Entanglement
where the von Neumann entanglement entropy
$S_{1}(\rho_{\mathrm{R}})=-\mathrm{Tr}(\rho_{\mathrm{R}}\log(\rho_{\mathrm{R}}))=-\sum\limits_{i}\lambda_{i}\log{\lambda_{i}},$ (7)
is recovered in the limit $\alpha\to 1^{+}$ and where $\lambda_{i}$ are eigenvalues of $\rho_{R}$. For random quadratic Hamiltonians, a class of models to which the SYK-2 Hamiltonian belongs [16], the average entanglement entropy was derived in closed form in [60], reading
$\mathcal{S}_{1}^{\mathrm{SYK-2}}\left(R,f\right)=\mathcal{K}(f)\ln{(2)}R,$ (8)
@@
Figure 2: Upper panels: Averaged bipartite von Neumann entanglement entropy, defined in Eq. (7), for the SYK-4 + SYK-2 model as a function of the interpolation parameter $g$. Shaded regions indicate the standard deviation over $M$ disorder realizations. Lower panels: Finite-size scaling of the relative entanglement entropy gap, defined in Eq. (11), along with its linear extrapolation to the thermodynamic limit. This figure highlights the fundamental difference in entanglement (i.e., non-local correlations) between the ground state (GS) and a middle-spectrum eigenstate (MS) across the interpolation. The left column shows results for the GS, while the right column corresponds to a MS.
Here, $f=R/N^{\rm spin}$ represents the ratio between the subsystem size $R$ and the total system size in terms of the number of qubits. The entanglement entropy scaling of the SYK-2 eigenstates follows a volume law (extensive scaling with subsystem size) but with a coefficient dependent on the subsystem-to-system size ratio $f$, distinguishing it from the fully quantum chaotic regime in the thermodynamic limit. Efforts to extend these insights to the SYK-4 model include works such as Refs. [61, 62, 17].
The maximally chaotic random pure state, sampled uniformly according to the Haar measure, exhibits the Page value for entanglement entropy, given by
@@
In upper row panels of Fig. 2, we show the rescaled average entanglement entropy in the ground state (GS) and middle-of-the-spectrum (MS) eigenstates of the interpolated SYK Hamiltonian $H(g)$, with $f=1/2$. The average is taken over subsystem bipartitions (denoted $\overline{S}$) and disorder realizations (denoted $\langle\cdot\rangle_{M}$) for each value of the interpolation parameter $g$. Disorder statistics for various system sizes are detailed in Appendix C.
[…]
For $g=0$ (SYK-4), the GS entanglement entropy approaches the Haar (universal) value more closely than at $g=1$ (SYK-2), though neither reaches it. This is quantified by the relative gap
$\Delta S_{1}:=\left|\dfrac{S_{1}-S_{1}^{\rm Haar}}{S_{1}^{\rm Haar}}\right|,$ (11)
whose finite-size scaling is shown in the lower panels. In the lower left panel of Fig. 2, the blue and orange dashed lines represent linear extrapolations to the thermodynamic limit $1/N\to 0$, where we observe a finite, non-zero value. In the lower panels we fitted the data to the function of the form $f(x)=a+bx$. For the $g=0$ case, the extrapolated values are $a=0.1\pm 0.02$ (GS) and $a=0.007\pm 0.017$ (MS), indicating a vanishing entanglement gap in the thermodynamic limit for the highly chaotic regime. In contrast, the MS entanglement shows that the SYK-4 model achieves near-perfect agreement with the Haar value, while SYK-2 does not. The behavior across $g$ indicates that universal features of MS states persist for all $g\neq 1$, highlighting the robustness of SYK-4 features in the higher energy states. To summarize, while it is well-known that the entanglement in SYK-4 obeys volume law for all the states, including the GS, we find that the GS deviates from the Haar value, while the MS states actually do reach the universal Haar value. This is not in contradiction with [17] as we compute entanglement exactly at half bipartition.
III.2 Entanglement Spectrum
A finer probe into the structure of entanglement is given by the full distribution of the eigenvalues of the reduced density Matrix (RDM) which we denote $\eta(x)$. For Haar-random pure states, this distribution corresponds to the Marchenko–Pastur (M-P) law, which arises as the limiting eigenvalue distribution of Wishart matrices [41, 65], and is given by
$\eta^{\rm Haar}(x)=1-\frac{2}{\pi}\left(x\sqrt{1-x^{2}}+\arcsin{x}\right),$ (12)
To distinguish two probability distributions we employ the Kullback-Leibler divergence
$D_{KL}(p||q):=\sum_{i}p_{i}(\log p_{i}-\log q_{i}).$ (13)
[…]
Figure 3: Normalized reduced density matrix (RDM) eigenvalues of $f=1/2$ subsystem-to-system. These are average values over many iterations C across a single bipartition. One obtains similar results choosing other bipartitions. The reference value is the Marchenko-Pastur (M-P) distribution $\eta^{\rm Haar}(x)$ in Eq. (12). We define $\eta_{k}=k/d$ and $x_{k}=(1/2)\sqrt{\lambda_{k}d}$ where $d=2^{N/4}$ for $f=1/2$, while the binning step for the interpolation parameter has been set to $\epsilon=0.01$. The left column shows results for the ground state (GS), while the right column corresponds to a middle-of-spectrum (MS) eigenstate.
In the left panel of Fig. 4, we show how the KL divergence–based fidelity behaves under different choices of the interpolation binning parameter $\epsilon$, for a fixed number of Majorana fermions. We find that the rescaled fidelity diagnostic introduced in this work exhibits a consistent response across various binning choices. As a byproduct, this robustness provides justification for the specific choice of $\epsilon=0.01$ used in Fig. 3. Moreover, the stability of the response allows us to reliably use this binning to extract the location of the suspected transition point. To do so, we fit the data to a high-degree polynomial function, enabling a binning-independent estimate of the transition point (shown as the solid black curve). We found that a degree-10 polynomial captures the relevant feature, i.e. the location of the peak sufficiently well. The peak position, indicated by the dashed vertical line, marks the critical point for the transition.
In the right panel of Fig. 4, we plot the transition point extracted from the polynomial fits as a function of system size $N$. The resulting scaling behavior is well-described by the power law
$g_{c}=N^{-0.78\pm 0.03}\sim N^{-3/4}.$ (14)
To the best of our knowledge, this is the first systematic numerical study that identifies this subtle crossover in the GS of the perturbed SYK model. Our analysis shows that the non-Fermi liquid GS of SYK-4 is unstable under SYK-2 type quadratic perturbations. Physically, this implies that the transition to Fermi liquid behavior occurs at a perturbation strength that vanishes in the thermodynamic limit, rendering the special SYK-4 character of the GS unstable for arbitrarily small quadratic terms as $N\to\infty$.
Figure 4: Left panel: KL divergence–based fidelity as a function of the interpolation parameter $g$ for different binning choices $\epsilon$, with fixed system size $N=20$. The rescaled fidelity diagnostic exhibits robust behavior across $\epsilon$, justifying the use of $\epsilon=0.01$ to extract the transition point. The solid black line represents a degree-10 polynomial fit used to locate the peak (marked by the dashed vertical line). Right panel: Finite-size scaling of the extracted transition point $g_{c}$, obtained from the location of the fidelity maximum. The fit reveals a power-law scaling, indicating that the SYK-4 ground state becomes unstable under arbitrarily small quadratic perturbations in the thermodynamic limit. In between $N^{-1}$ and $N^{-1/2}$ is where Ref. [14] identified the potential transition to be located.
III.3 Entanglement Spectrum Statistics (ESS)